# Existence and Uniqueness for a Porous-Medium-Type Diffusion Model with Time-Dependent Growth

### v2

@@ 2. Proof of the Main Results
In the following proof of the Theorem 1.1, we apply the successive approximation method, as our goal is to implement the result using a computer language.
@@
Then, the function $\underline{u}(x)=\sigma\phi_{1}$ is a subsolution of (1.4) while $\overline{u}\left(x\right)=cw(x)$ is a supersolution. Indeed,
@@
Next, we define the sequence $\{u^{k}\}_{k\geq 0}$ by
@@
Using the Brezis-Oswald inequality [3, Theorems 1 and 2], we derive
@@
Since the sequence $\{u^{k}\}_{k\geq 0}$ is monotone increasing and bounded above (by $\overline{u}(x)$), it follows by the monotone convergence theorem that the pointwise limit
@@
The uniqueness of the solution to the problem (1.4) follow again from the results established by Brezis and Oswald [3, Theorems 1 and 2] (see also [1] or [2]).
Now, let us prove the validity of (1.5). Assume a solution of the form
@@
We proceed by establishing the proof of the comparison principle.
@@
Let $\mathcal{L}_{\mu}[u]:=\frac{\partial u}{\partial t}-\Delta u^{\alpha}-\mu(t)u$ be the parabolic operator associated with (1.1). By hypothesis, $v$ and $w$ are the unique classical solutions satisfying $\mathcal{L}_{\mu_{1}}[v]=0$ and $\mathcal{L}_{\mu_{2}}[w]=0$ in $\Omega\times(0,T)$. We first establish that $v$ serves as a subsolution for the operator $\mathcal{L}_{\mu_{2}}$. Indeed, substituting $v$ into $\mathcal{L}_{\mu_{2}}$ yields:
@@
Since $\mu_{1}(t)\leq\mu_{2}(t)$ for all $t$ and $v(x,t)\geq 0$ in $\Omega\times(0,T)$, it follows that $(\mu_{1}(t)-\mu_{2}(t))v\leq 0$, which implies:
@@
Furthermore, the initial and boundary conditions satisfy $v(x,0)=v_{0}(x)\leq w_{0}(x)=w(x,0)$ in $\Omega$ and $v(x,t)=0=w(x,t)$ on $\partial\Omega\times(0,T)$.
To establish the ordering $v\leq w$, we define z(x,t):=v(x,t)w(x,t)z(x,t):=v(x,t)-w(x,t) and note that $z$ satisfies:
@@
Using the monotonicity of the mapping $s\mapsto s^{\alpha}$ for $\alpha>1$ and $s\geq 0$, we can express the difference $v^{\alpha}-w^{\alpha}$ as c(x,t)(vw)c(x,t)(v-w) where $c(x,t)=\int_{0}^{1}\alpha[w+\theta(v-w)]^{\alpha-1}d\theta\geq 0$. Substituting this into the inequality for $z$, we obtain:
@@
Given that $z(x,0)\leq 0$ and $z$ vanishes on the boundary, the application of the maximum principle for porous-medium-type equations (see [14, Chapter 3] or [11, Theorem 2.1]) ensures that $z(x,t)\leq 0$ on $\overline{\Omega}\times[0,T)$. Consequently, $v(x,t)\leq w(x,t)$ for all $(x,t)\in\overline{\Omega}\times[0,T)$, which completes the proof. ∎
We end the section by establishing the proof of the third main result.
@@
We first compute the quantities appearing in the parabolic operator for the functions $v_{\pm}$.
@@
which shows that $v_{-}$ is a subsolution of (1.1). On the boundary, since $u_{-}=0$ on $\partial\Omega$, we have
@@
(b) Supersolution case. Assume $u=u_{+}$ satisfies (1.9), i.e.
@@
and $v_{+}$ is a supersolution of (1.1). The boundary condition $v_{+}=0$ on $\partial\Omega\times(0,T)$ follows from $u_{+}=0$ on $\partial\Omega$.
@@
$v_{-}(x,t)=S(t)\,u_{-}^{1/\alpha}(x),\qquad v_{+}(x,t)=S(t)\,u_{+}^{1/\alpha}(x),\quad(x,t)\in\overline{\Omega}\times[0,T),$
so the initial data is sandwiched between the sub- and supersolution at the time $t=0$. Moreover, for all $x\in\overline{\Omega}$ yields and $t\in[0,T)$ we have
@@
To complete the argument, we now invoke the classical subsolution–supersolution theory for nonlinear parabolic equations. We proceed in several steps.
@@
Since $v_{-}(x,0)=\gamma\,u_{-}^{1/\alpha}(x)$ and $v_{+}(x,0)=\gamma\,u_{+}^{1/\alpha}(x)$, and because $u_{-}\leq u_{+}$ in $\overline{\Omega}$, this ordering is consistent with the assumptions.
@@
Step 3.3: Application of comparison principles. The theory developed by Pao [11] (see also [8] for related results) establishes that, under suitable regularity assumptions, if $v_{-}$ and $v_{+}$ are respectively an ordered subsolution and supersolution of the parabolic problem (1.1) with boundary conditions (1.2), and if they satisfy the initial ordering at $t=0$, then there exists a unique classical solution $v\in C^{2,1}(\Omega\times(0,T))\cap C(\overline{\Omega}\times[0,T))$ of the problem (1.1)–(1.2)–(1.3) such that
@@
Step 3.4: Verification of regularity hypotheses. The regularity assumptions required by the theory are satisfied:
@@
is well-defined and that the subsolution–supersolution method applies.
@@
(a) Choice of monotonicity constant. We first choose a function $c(x,t)\geq 0$ sufficiently large to ensure monotonicity. Specifically, consider the nonlinearity in (1.1):
@@
(b) Iterative scheme. Define two sequences $\{v^{(m)}_{-}\}_{m=0}^{\infty}$ and $\{v^{(m)}_{+}\}_{m=0}^{\infty}$ as follows. Initialize with
$v^{(0)}_{-}(x,t):=v_{-}(x,t),\qquad v^{(0)}_{+}(x,t):=v_{+}(x,t).$ (2.18)
For $m\geq 1$, construct $v^{(m)}_{-}$ and $v^{(m)}_{+}$ by solving the linearized parabolic problem
@@
(c) Monotonicity and convergence. By the comparison principle for parabolic equations and the choice of $c(t)$, the sequences satisfy
$v_{-}=v^{(0)}_{-}\leq v^{(1)}_{-}\leq v^{(2)}_{-}\leq\cdots\leq v^{(m)}_{-}\leq\cdots\leq v^{(m)}_{+}\leq\cdots\leq v^{(2)}_{+}\leq v^{(1)}_{+}\leq v^{(0)}_{+}=v_{+}.$ (2.20)
That is, the lower sequence $\{v^{(m)}_{-}\}$ is monotone increasing and bounded above by $v_{+}$, while the upper sequence $\{v^{(m)}_{+}\}$ is monotone decreasing and bounded below by $v_{-}$. Standard parabolic regularity estimates (see [11], Theorem 2.1) ensure that both sequences converge uniformly on compact subsets of $\overline{\Omega}\times[0,T)$ to limits
$\underline{v}(x,t):=\lim_{m\to\infty}v^{(m)}_{-}(x,t),\qquad\overline{v}(x,t):=\lim_{m\to\infty}v^{(m)}_{+}(x,t).$
Moreover, both $\underline{v}$ and $\overline{v}$ are classical solutions of (1.1)–(1.2)–(1.3), and they are the minimal and maximal solutions, respectively, in the ordered interval $v_{-},v_{+}]$.
@@
This constructive approach is particularly well-suited for numerical implementation: at each iteration $m$, one solves a linear parabolic problem (since the right-hand side depends only on the previous iterate $v^{(m-1)}$), which can be discretized using standard finite difference or finite element methods. The monotone convergence (2.20) provides a natural stopping criterion based on the $L^{\infty}$ or $L^{2}$ norm of the difference $\|v^{(m)}-v^{(m-1)}\|$.
Step 3.5: Uniqueness. Uniqueness of the classical solution follows from the monotonicity properties of the operator. Specifically, if $v^{(1)}$ and $v^{(2)}$ are two solutions both lying in the interval $v_{-},v_{+}]$, then the difference $w:=v^{(1)}-v^{(2)}$ satisfies
@@
By mean-value expansions of the nonlinear term and the maximum principle for parabolic equations, one concludes that $w\equiv 0$, yielding uniqueness.
Step 3.6: Conclusion. We have shown that $v_{-}$ and $v_{+}$ are respectively a subsolution and a supersolution of (1.1) with boundary condition (1.2), that they are ordered, and that the initial data lies between them. Invoking the standard subsolution–supersolution existence theorem (see [11], Theorem 1.3 and related results, or [8], Section 3), we conclude that there exists a unique classical solution $v$ of the problem (1.1)–(1.2)–(1.3) satisfying the pointwise estimate (2.16). This completes the proof. ∎
@@ 3. Derivation of the Model via Conservation Laws
where $v(x,t)$ represents the population density, $\mathbf{J}$ is the population flux, and $f(v,t)$ is the net growth rate. We consider a constitutive function
@@
In our model, the population flux $\mathbf{J}$ is governed by this potential $\varphi$, following a nonlinear Fick’s law:
@@
This indicates that the dispersal is driven by the gradient of the potential $\varphi(v)$. Substituting (3.2) into (3.1) yields the diffusion term $\nabla\cdot(\varphi^{\prime}(v)\nabla v)$.
This flux-driven dynamic can be intimately connected to a variational principle. The Euler–Lagrange conditions associated with the functional are inspired by the population flux $\mathbf{J}$, suggesting that the system evolves to minimize an associated energy. We consider:

### v3

@@ 3. Proof of the Main Results
In the following proof of Theorem 1.1, we apply the successive approximation method, because our goal is to implement the result using a computer language.
@@
Then, the function $\underline{u}(x)=\sigma\phi_{1}$ is a subsolution of (1.5) whereas $\overline{u}\left(x\right)=cw(x)$ is a supersolution. Indeed,
@@
Next, we define the sequence $\{u^{k}\}_{k\geq 0}$ as
@@
Using the Brezis-Oswald inequality [5, Theorems 1 and 2], we derive
@@
Because the sequence $\{u^{k}\}_{k\geq 0}$ is monotonically increasing and bounded above (by $\overline{u}(x)$), it follows from the monotone convergence theorem that the pointwise limit
@@
The uniqueness of the solution to problem (1.5) follows again from the results established by Brezis and Oswald [5, Theorems 1 and 2] (see also [1] or [4]).
Now, let us prove the validity of (1.4). Assume that a solution of the form
@@
Remark 3.1.
$v(x,t)=S(t)\cdot u^{1/\alpha}(x),$
as established in Theorem 1.1, is publicly available at [8].
Remark 3.2.
The proof of Theorem 1.1 reveals the essential mechanism behind the Bernoulli localization principle: the substitution of the product ansatz (3.4) into the parabolic PDE (1.1) automatically produces the Bernoulli ODE (3.7), because the spatial dependence cancels exactly when $u$ solves the elliptic problem (1.5). This cancellation is not an accident; it is a structural feature of the power-law nonlinearity $v^{\alpha}$, which ensures that $\Delta v^{\alpha}=S^{\alpha}\Delta u$ is proportional to $v$ through the elliptic equation. This structural observation is the conceptual heart of the localization principle.
We then establish a proof of the comparison principle.
@@
Let
$\mathcal{L}_{\mu}[u]:=\frac{\partial u}{\partial t}-\Delta u^{\alpha}-\mu(t)u$
be the parabolic operator associated with (1.1). By hypothesis, $v$ and $w$ are unique classical solutions satisfying
$\mathcal{L}_{\mu_{1}}[v]=0\text{ and }\mathcal{L}_{\mu_{2}}[w]=0\text{ in }\Omega\times(0,T).$
We first establish that $v$ serves as a subsolution for the operator $\mathcal{L}_{\mu_{2}}$. Indeed, substituting $v$ into $\mathcal{L}_{\mu_{2}}$ yields:
@@
Because $\mu_{1}(t)\leq\mu_{2}(t)$ for all $t$ and $v(x,t)\geq 0$ in $\Omega\times(0,T)$, it follows that $(\mu_{1}(t)-\mu_{2}(t))v\leq 0$, which implies:
@@
Furthermore, the initial and boundary conditions satisfy
$v(x,0)=v_{0}(x)\leq w_{0}(x)=w(x,0)\text{ in }\Omega\text{,}$
and
$v(x,t)=0=w(x,t)\text{ on }\partial\Omega\times(0,T).$
To establish the ordering $v\leq w$, we define
z(x,t):=v(x,t)w(x,t)z(x,t):=v(x,t)-w(x,t)
and note that $z$ satisfies:
@@
Using the monotonicity of the mapping $s\mapsto s^{\alpha}$ for $\alpha>1$ and $s\geq 0$, we can express the difference $v^{\alpha}-w^{\alpha}$ as c(x,t)(vw)c(x,t)(v-w) where
$c(x,t)=\int_{0}^{1}\alpha[w+\theta(v-w)]^{\alpha-1}d\theta\geq 0.$
Substituting this into the inequality for $z$, we obtain:
@@
Given that $z(x,0)\leq 0$ and $z$ vanishes at the boundary, the application of the maximum principle for porous-medium-type equations (see [24, Chapter 3] or [20, Theorem 2.1]) ensures that $z(x,t)\leq 0$ on $\overline{\Omega}\times[0,T)$. Consequently, $v(x,t)\leq w(x,t)$ for all $(x,t)\in\overline{\Omega}\times[0,T)$, completing the proof. ∎
We conclude this section by establishing the proof of the third main result.
@@
First we compute the quantities appearing in the parabolic operator for the functions $v_{\pm}$.
@@
which shows that $v_{-}$ is a subsolution of (1.1). On the boundary, because $u_{-}=0$ on $\partial\Omega$, we have
@@
(b) Supersolution case. Assume $u=u_{+}$ satisfies (1.9), that is
@@
where $v_{+}$ is the supersolution to (1.1). The boundary condition $v_{+}=0$ on $\partial\Omega\times(0,T)$ follows from $u_{+}=0$ on $\partial\Omega$.
@@
$v_{-}(x,t)=S(t)\,u_{-}^{1/\alpha}(x),\qquad v_{+}(x,t)=S(t)\,u_{+}^{1/\alpha}(x),\quad(x,t)\in\overline{\Omega}\times[0,T).$
Therefore the initial data are sandwiched between the sub- and supersolutions at time $t=0$. Moreover, for all $x\in\overline{\Omega}$ yields and $t\in[0,T)$ we have
@@
To complete this argument, we invoke classical subsolution–supersolution theory for nonlinear parabolic equations. We proceeded in several steps.
@@
Since
[…]
and because $u_{-}\leq u_{+}$ in $\overline{\Omega}$, this ordering is consistent with the assumptions.
@@
Step 3.3: Application of comparison principles. The theory developed by Pao [20] (see also [16] for related results) establishes that, under suitable regularity assumptions, if $v_{-}$ and $v_{+}$ are respectively an ordered subsolution and supersolution of the parabolic problem (1.1) with boundary conditions (1.2), and if they satisfy the initial ordering at $t=0$, then there exists a unique classical solution
$v\in C^{2,1}(\Omega\times(0,T))\cap C(\overline{\Omega}\times[0,T)),$
of the problem (1.1)–(1.2)–(1.3) such that
@@
Step 3.4: Verification of the regularity hypotheses. The regularity assumptions required by the theory are satisfied:
@@
is well defined and the subsolution–supersolution method applies.
@@
(a) Choice of monotonicity constant. We first choose a function $c(x,t)\geq 0$ that is sufficiently large to ensure monotonicity. Specifically, we consider the nonlinearity in (1.1), as follows:
@@
(b) Iterative scheme. We define the two sequences $\{v_{-}^{(m)}\}_{m=0}^{\infty}$ and $\{v_{+}^{(m)}\}_{m=0}^{\infty}$ as follows. Initialize with
$v_{-}^{(0)}(x,t):=v_{-}(x,t),\qquad v_{+}^{(0)}(x,t):=v_{+}(x,t).$ (3.18)
For $m\geq 1$, construct $v_{-}^{(m)}$ and $v_{+}^{(m)}$ by solving the linearized parabolic problem
@@
(c) Monotonicity and convergence. According to the comparison principle for parabolic equations and the choice of $c(t)$, the sequences satisfy
$v_{-}=v_{-}^{(0)}\leq v_{-}^{(1)}\leq v_{-}^{(2)}\leq\cdots\leq v_{-}^{(m)}\leq\cdots\leq v_{+}^{(m)}\leq\cdots\leq v_{+}^{(2)}\leq v_{+}^{(1)}\leq v_{+}^{(0)}=v_{+}.$ (3.20)
That is, the lower sequence $\{v_{-}^{(m)}\}$ monotonically increases and is bounded above by $v_{+}$, whereas the upper sequence $\{v_{+}^{(m)}\}$ monotonically decreases and is bounded below by $v_{-}$. Standard parabolic regularity estimates (see [20], Theorem 2.1) ensure that both sequences converge uniformly on compact subsets of $\overline{\Omega}\times[0,T)$ to limits
$\underline{v}(x,t):=\lim_{m\rightarrow\infty}v_{-}^{(m)}(x,t),\qquad\overline{v}(x,t):=\lim_{m\rightarrow\infty}v_{+}^{(m)}(x,t).$
Moreover, both $\underline{v}$ and $\overline{v}$ are classical solutions of (1.1)–(1.2)–(1.3), and they are the minimal and maximal solutions, in the ordered interval $v_{-},v_{+}]$.
@@
This constructive approach is particularly well-suited for numerical implementation. At each iteration $m$, one solves a linear parabolic problem (because the right-hand side depends only on the previous iterate $v^{(m-1)}$), which can be discretized using standard finite difference or finite element methods. Monotone convergence (3.20) provides a natural stopping criterion based on the $L^{\infty}$ or $L^{2}$ norm of the difference $\|v^{(m)}-v^{(m-1)}\|$.
Step 3.5: Uniqueness. The uniqueness of the classical solution is derived from the monotonicity of the operator. Specifically, if $v^{(1)}$ and $v^{(2)}$ are two solutions that lie in the interval $v_{-},v_{+}]$, then the difference $w:=v^{(1)}-v^{(2)}$ satisfies
@@
By mean-value expansions of the nonlinear term and the maximum principle for parabolic equations, we conclude that $w\equiv 0$, yielding uniqueness.
Step 3.6: Conclusion. We have shown that $v_{-}$ and $v_{+}$ are respectively a subsolution and a supersolution of (1.1) with boundary condition (1.2), that they are ordered, and that the initial data lie between them. Invoking the standard subsolution–supersolution existence theorem (see [20], Theorem 1.3 and related results, or [16], Section 3), we conclude that there exists a unique classical solution $v$ of problem (1.1)–(1.2)–(1.3) satisfying the pointwise estimate (3.16). This completes this proof. ∎
Remark 3.3.
The Python code available at: [9] implements Theorem 1.3 using the subsolution–supersolution framework. In the special case where the solution admits a separable form
$v(x,t)=S(t)\cdot u^{1/\alpha}(x),$
Remark 3.4.
The Python script available at: [10] extends the implementation to illustrate Theorem 1.3 under the general initial condition proposed in Remark 1.4. In particular, the parabolic solver is initialized with a weighted average of the sub- and supersolutions,
$v_{0}(x)=\frac{7}{8}v_{-}(x,0)+\frac{1}{8}v_{+}(x,0),$
thereby demonstrating that the monotone iteration scheme robustly captures the existence and uniqueness of the solutions for arbitrary initial data within the ordered interval $v_{-},v_{+}]$.
@@ 4. Derivation of the Model via Conservation Laws
where $v(x,t)$ is the population density, $\mathbf{J}$ is the population flux, and $f(v,t)$ is the net growth rate. The following constitutive function is considered:
@@
In our model, the population flux $\mathbf{J}$ is governed by the potential $\varphi$, following a nonlinear Fick’s law:
@@
This indicates that dispersal is driven by the gradient of the potential $\varphi(v)$. Substituting (4.2) into (4.1) yields the diffusion term $\nabla\cdot(\varphi^{\prime}(v)\nabla v)$.
This flux-driven dynamic can be connected to a variational principle. The Euler–Lagrange conditions associated with the functional are inspired by the population flux $\mathbf{J}$, which suggest that the system evolves to minimize the associated energy. We consider: